# FineFake: A Knowledge-Enriched Dataset for Fine-Grained Multi-Domain Fake News Detection

### v2

@@ Abstract.
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06
1. Introduction
In the contemporary landscape of the ever-evolving digital society, social media stands as a prominent medium for accessing news. It has emerged as an optimal milieu for the dissemination of falsified information, posing a significant threat to both individuals and society (zhou2020survey, ; zhang2024reinforced, ; li2017ppne, ; zhao2022learning, ). For example, within the COVID-19 infodemic, the propagation of counterfeit news has engendered social unrest and precipitated a multitude of fatalities stemming from erroneous medical interventions (rocha2021impact, ; greene2021quantifying, ; yan2023comprehensive, ; zhang2023efficiently, ). Hence, the automatic identification of spurious news has recently assumed paramount importance, thereby eliciting substantial scholarly dedication. Alongside the array of advanced detection models, a comprehensive and precise benchmark dataset is imperative for the efficacy of fake news detection. Such a dataset enables a thorough reflection of the multifaceted nature and intricacies inherent in fake news, thereby furnishing a more precise reference for both model training and evaluation. In order to enhance the pursuit of identifying fake news, a series of datasets have been put forth as evidenced in Table 2. Commencing with small scale and unimodal ones, the accessible datasets have progressed into expansive-scale compilations, incorporating multiple modalities. Such advancements have significantly augmented the wealth of extensive and diverse information available for further scrutiny and analysis.
Table 1. Performance (accuracy) of MVAE under the cross-platform and cross-topic settings. The abbreviations of topics are: Pol: Politics; Ent.: Entertainment; Con.: Conflict.
@@ 1. Introduction
Notwithstanding the notable advancements in fake news detection datasets, existing datasets are generally constructed upon the news centered around a similar topic or a single platform, leading to the limited generality. For instance, as illustrated in Table 2, the LIAR and Breaking datasets (wang2017liar, ; pathak2019breaking, ) only comprise samples related to the topic of politics, whereas the Weibo and Twitter datasets (jin2017multimodal, ; boididou2015verifying, ) solely consist of news sourced from a single platform. Nevertheless, within the sphere of different real-world news platforms, an incessant deluge of millions of news articles spanning various topics. The diverse range of topics and platforms is largely ignored by existing datasets, leading to inadequate assessments of cross-domain capability. For example, as depicted in Table 1, when a widely-used detection model MVAE (khattar2019mvae, ) is trained on a specific news topic or platform and then applied to another topic or platform, its performance exhibits a notable decline. The underlying reasons may be attributed to two key aspects. From the perspective of semantic topics, tokens commonly associated with fake news within certain domains, such as “vaccine”, “virus”, and “side-effect” in the health domain, may be notably scarce in other domains like business. Consequently, the semantic distributions of news across topics are apt to diverge significantly, introducing the classical covariate shift problem (shimodaira2000improving, ). From the perspective of platforms, the proportion of fake news can vary significantly across different platforms. For instance, reputable sources like CNN generally feature higher credibility compared to self-media posts on Twitter, which may have a higher proportion of true news (bovet2019influence, ). This imbalance in the proportion of real and fake news across platforms introduces another classic challenge in domain adaptation: the label shift problem (garg2020unified, ). These two domain shifts underscore the necessity for a benchmark dataset for fake news to evaluate a model’s domain adaptation capabilities.
To alleviate the aforementioned challenges of domain shift, external knowledge graphs like ConceptNet (speer2017conceptnet, ) are incorporated as complementary to provide extra cross-domain information, serving as a bridge between disparate domains. In the area of fake news detection, prior research endeavors (dun2021kan, ; chen2022evidencenet, ; hu2021compare, ) have also substantiated the significance of leveraging common knowledge derived from external knowledge sources (e.g., wikidata) to enhance the precision of fake news detection. The determination of the veracity of news often relies on foundational knowledge and latest external information as auxiliary factors for assessment. Hence, accurate supplementary knowledge plays a crucial role in cross-domain fake news detection tasks. However, the majority of existing datasets do not include associated knowledge or rely on evidences sourced from online websites, introducing potential inaccuracies and extraneous noise information. Furthermore, it is common for the same entity to possess varied referential meanings across different domains. As depicted in Fig. 1, the term “apple” within the health domain signifies a type of fruit, whereas in the business domain, it denotes the company Apple Inc. Consequently, these disparities across domains give rise to the challenge of entity ambiguity, which can introduce noisy knowledge and degrade the performance of the model. Therefore, the development of datasets that encompass accurate common knowledge assumes an imperative role in advancing the progress of fake news detection area.
@@
Figure 1. An illustration of FineFake dataset, which is a multi-domain dataset that encompasses instances collected from diverse platforms and topics. Each sample is associated with accurate knowledge and fine-grained label.
Another significant aspect of fake news across different topics or platforms lies in the diverse underlying strategies used for fabrication. For instance, image-based fake news comprises the highest proportion, around 40%, among fake news on Reddit, whereas fake news from CNN contains less than 5% in the image-based fake category within the FineFake dataset. This discrepancy can be attributed to the higher prevalence of manipulated images on social media platforms compared to official news sources (nygren2018changing, ). Therefore, it is essential to identify distinguishing features of fake news that can provide a rational basis for judgment across platforms. However, conventional fake news detection datasets typically classify news articles into binary classes (true or fake) or employ broad categories such as “most likely” or “somewhat likely”. Unfortunately, such coarse-grained annotations fail to elucidate the underlying factors contributing to the falseness of a news item, such as fabricated images or inconsistencies between text and image. Such capabilities are crucial for achieving a comprehensive and deep understanding of the intrinsic reasons behind fake news. Hence, there is an urgent need for the development of a fine-grained annotation strategy capable of unraveling the underlying reasons behind fake news.
@@
Table 2. Comparison between FineFake and other fake news detection datasets.
@@
LIAR(wang2017liar, ) 12,836 Politifact Political Six Category
Breaking!(pathak2019breaking, ) 649 BS Detector US Election Three Category
Weibo21(nan2021mdfend, ) 9128 Weibo Nine Topics Real/ Fake
Evons(krstovski2022evons, ) 92,969 Media-source US Election Real/ Fake
Weibo(jin2017multimodal, ) 9,528 Weibo Real/ Fake
Pheme(zubiaga2017exploiting, ) 5,802 Twitter News Events Real/ Fake
FauxBuster(zhang2018fauxbuster, ) 917 Twitter/Reddit Real/ Fake
Twitter(boididou2015verifying, ) 15,629 Twitter Real/ Fake
MM-Covid(li2020mm, ) 11,173 FullFact Health Real/ Fake
MuMIN(nielsen2022mumin, ) 984 Twitter Three Category
$MR^{2}$(hu2023mr2, ) 14,700 Twitter/Weibo Three Category
@@
FineFake 16,909 Snopes, Twitter, Reddit CNN, Apnews, Cdc.gov Nytimes, Washingtonpost Politics, Enter. Business, Health Society, Conflict Real, Text/ Image Fake, Text-image/ Content-knowledge Inconsistency, Others
@@
In order to address the aforementioned challenges, in this paper we propose a comprehensive and knowledge-enhanced dataset for fake news detection, dubbed FineFake. As illustrated in Table 2, FineFake surpasses its predecessors by spanning multiple topics and platforms, enjoying accurate common knowledge and fine-grained annotations, thereby furnishing robust and solid data support for further research endeavors. FineFake encompasses instances collected from diverse media platforms, such as CNN, Reddit and Snopes. All instances are also labeled into six distinct topics (i.e., politics, entertainment, business, health, society and conflict). The inclusion of this comprehensive multi-domain dataset fosters a deeper comprehension of correlations between fake news from various domains. Each news article contains textual content, images, possible social connections, and other pertinent meta-data. To ensure the provision of reliable common knowledge, each news is appended with the relevant knowledge entities and descriptions in a semi-manual labeling manner. Moreover, we introduce an innovative annotation guideline that extends beyond traditional binary class labels. Our approach incorporates a six-category annotation strategy that sheds light on the reasons behind the detected fake news, including real, textual fake, visual fake, text-image inconsistency, content-knowledge inconsistency, and other samples. Based on FineFake, we conduct extensive experiments, including investigations into data characteristics, domain-adaptation performance analysis, and exploration of fine-grained classification results, thus providing a highly desirable benchmark for future investigations in this domain. Furthermore, to solve the covariate shift problem and label shift problem in FineFake, we propose a knowledge-enhanced domain adaptation network, dubbed KEAN. Extensive experiments demonstrate that our method achieves SOTA performance across the majority of scenarios. Our contributions are summarized as:
@@
  • To the best of our knowledge, FineFake dataset represents a pioneering effort for fake news detection. It systematically gathers and formalizes multi-modal news content from diverse topics and platforms, presenting a valuable resource for advancing research under cross-domain scenarios.
@@
2. Related Work
2.1. Fake News Detection Datasets
@@
Since the rapid development of the internet, a proliferation of publicly available datasets concerning the detection of fake news has ensued. Initially, researchers predominantly focused on collecting textual data, concentrating on specific domains to construct these datasets (wang2017liar, ; thorne2018fever, ; pathak2019breaking, ; patwa2021fighting, ). For instance, the LIAR dataset (wang2017liar, ) harnesses Politifact [^1] http://www.politifact.com/ to extract news items, subsequently annotating them with six classification labels. Similarly, FEVER (thorne2018fever, ) is generated by altering sentences extracted from Wikipedia and pre-processing Wikipedia data, each claim is annotated to a three-way classification label. Additionally, the COVID-19 dataset (patwa2021fighting, ) meticulously curates a manually annotated repository of social media posts and articles pertaining to COVID-19, aiming to facilitate research endeavors in identifying pertinent rumors that possess the potential to instigate significant harm.
@@
The aforementioned uni-modal datasets primarily concentrate on linguistic analysis, thus disregarding the crucial dimensions of social networks for dissemination, corresponding images, and metadata, essential for a more comprehensive detection framework. In stark contrast to traditional textual news media releases, the presence of multimodal news incorporating images or videos tends to captivate greater attention and propagate more extensively (zhang2023cisum, ), thus may lead to more damage transmission. Consequently, there has been a discernible surge in the construction of datasets integrating images and social network data (shu2020fakenewsnet, ; zhang2018fauxbuster, ; nielsen2022mumin, ; zubiaga2017exploiting, ; boididou2015verifying, ). For instance, MM-COVID (li2020mm, ) offers a multilingual dataset encompassing news articles augmented with pertinent social context and images, aimed at facilitating the detection and mitigation of fake news pertaining to the COVID-19 pandemic. Similarly, Weibo (jin2017multimodal, ) collects original tweet texts, attached images, and contextual information sourced from Weibo, a prominent Chinese microblogging platform renowned for its objective ground-truth labels.
@@
Although the existing multi-modal datasets contain multi-modal data like images, the majority of them overlook critical factors such as multi-domain attributions, external knowledge and fine-grained classification annotations. $MR^{2}$ (hu2023mr2, ) attempts to address this gap by incorporating evidence retrieved from online sources as metadata to enhance fake news detection. Nonetheless, this approach of online retrieval lacks a guarantee of the accuracy of external evidence, potentially introducing extraneous noise information instead. Additionally, Weibo21  (nan2021mdfend, ) divides textual news data into nine distinct topics, yet it only focus on topics, overlooking the the multitude of platforms through which news dissemination occurs.
2.2. Fake News Detection Methods
@@
At the outset, a considerable amount of research focused on refining the extraction of semantic features inherent in news content itself, recognizing the wealth of information embedded within the content conducive to discerning its veracity (ma2016detecting, ; khattar2019mvae, ; Zhou2020SAFESM, ; chen2022cross, ). However, the escalating convergence of semantic structures between fake and authentic news has rendered the task of distinguishing between them based solely on semantics increasingly formidable (zhou2020survey, ). Consequently, attention has shifted towards leveraging external knowledge as supplementary information to bolster fake news detection efforts (qian2021knowledge, ; dun2021kan, ; sun2023inconsistent, ; sun2021inconsistency, ; hu2021compare, ). For instance, CompareNet (hu2021compare, ) constructs a directed heterogeneous document graph to compare news to external knowledge base through the extractied entities.
@@
Nonetheless, news content exhibits significant variations across diverse platforms and topics, and the distribution of fake and authentic news also fluctuates accordingly (zhou2020survey, ; zhu2022memory, ). Effective deployment of a well-trained fake news detection model in real-world scenarios necessitates robust cross-domain capabilities. However, only a limited number of studies have earnestly tackled the challenges posed by multi-domain and cross-domain fake news detection (wang2018eann, ; nan2021mdfend, ; zhu2022memory, ; yue2022contrastive, ). Consequently, we construct a multi-domain knowledge-enhanced multimodal fine-grained dataset that holds immense potential for facilitating research in the realms of multi-domain and cross-domain fake news detection in reality.
@@
3. FineFake: The Multi-Domain Dataset
This section presents an in-depth exploration of the fundamental phases involved in constructing the dataset, encompassing multi-domain news collection, semi-manual knowledge alignment, and fine-grained human annotation.
3.1. Multi-domain News Collection
Current fake news detection datasets predominantly concentrate on sourcing data from singular platforms within narrow topics. Nonetheless, this approach is encumbered by limitations pertaining to both data scale and diversity, thereby restricting the adaptability of detection models within varied real-world news landscapes. To overcome this constraint and develop a large-scale dataset spanning multiple platforms and topics with substantial diversity, we employ a comprehensive data collection strategy that encompasses three primary channels: fact-checking websites (e.g., snopes.com), social media platforms (e.g., Twitter), and reputable news websites (e.g., CNN). Through this multifaceted approach, we envisage the creation of a dataset that mirrors the complexity and heterogeneity inherent in contemporary news dissemination paradigms, thereby providing a benchmark to furnish researchers with a more robust foundation for advancing fake news detection methodologies.
@@ 3.1. Multi-domain News Collection
Table 3. Mappings from the categories in Snopes to topics.
@@
In this study, we leverage Snopes [^2] https://www.snopes.com/sitemap/ as a pivotal resource for acquiring verified news content categorized by various topics, leveraging the category labels meticulously assigned by professional editors. Snopes serves as a fact-checking website, furnishing invaluable research materials utilized by seasoned editors in their preparatory endeavors. Consequently, each verified claim encompasses a substantial amount of additional information regarding the original news post, including external links to the original source, expert-assigned topic labels, and authenticity assessments provided by professionals. The correlation between the category labels in Snopes and our designated topic labels is elucidated in Table 3. For each verified claim on Snopes, we meticulously compile all pertinent components sourced from the fact-checking websites, employing the external links furnished within each claim to procure news content from traditional official media sources and online streaming platforms.
Upon gathering all external links, we embark on a comprehensive data analysis endeavor to identify additional websites for data collection. For traditional official media sources, including APNews, CNN, New York Times, The Washington Post, and CDC333www.apnews.com, www.cnn.com, www.nytimes.com, www.washingtonpost.com, www.cdc.gov, we develop custom crawlers tailored to retrieve textual content, images, and associated metadata such as authors and publication dates for each news article.
In the realm of online streaming media, we leverage two prominent platforms: Twitter444www.twitter.com and Reddit555www.reddit.com. To collect user interactions associated with posted news and construct the social network, we utilize the Twitter API [^6] https://developer.twitter.com/en/docs/twitter-api/getting-started/about-twitter-api. Although Reddit lacks a designated API endpoint for retrieval, we engineer a specialized crawler to gather posts and construct the social network accordingly.

### v3

@@ Abstract
Index Terms:
Fake News Detection, Multi-Domain Benchmark, Fine-Grained Classification.
I Introduction
In the contemporary landscape of the ever-evolving digital society, social media stands as a prominent medium for accessing news. It has emerged as an optimal milieu for the dissemination of falsified information, posing a significant threat to both individuals and society [1]. For example, during the COVID-19 infodemic, the spread of fake news caused incorrect medical interventions, leading to social unrest and numerous fatalities [2, 3]. Hence, automatic fake news detection has become crucial, drawing significant academic focus. To enhance the pursuit of identifying fake news, a series of datasets like Twitter [4] and Pheme [5] has been developed, evolving from small, unimodal sets to large, multimodal compilations. These enhancements have augmented the wealth of extensive and diverse information available for further analysis.
TABLE I: Performance (accuracy) of MVAE under the cross-platform and cross-topic settings. The abbreviations of topics are: Pol: Politics; Ent.: Entertainment; Con.: Conflict.
@@ I Introduction
Notwithstanding the notable advancements in fake news detection datasets, existing datasets are generally constructed upon the news centered around a similar topic or a single platform, leading to the limited generality. For instance, as illustrated in Table II, the LIAR and Breaking datasets [6, 7] only comprise samples related to the topic of politics, whereas the Weibo and Twitter datasets [8, 4] solely consist of news sourced from a single platform. Nevertheless, within the sphere of different real-world news platforms, an incessant deluge of millions of news articles spanning various topics. The diverse range of topics and platforms is largely ignored by existing datasets, leading to inadequate assessments of cross-domain capability. For example, as depicted in Table I, when a widely-used detection model MVAE [9] is trained on a specific news topic or platform and then applied to another topic or platform, its performance exhibits a notable decline. The underlying reasons may be attributed to two key aspects. From the perspective of semantic topics, tokens commonly associated with fake news within certain domains, such as “vaccine”, “virus”, and “side-effect” in the health domain, may be notably scarce in other domains like business. Consequently, the semantic distributions of news across topics are apt to diverge significantly, introducing the classical covariate shift problem [10]. From the perspective of platforms, the proportion of fake news can vary significantly across different platforms. For instance, reputable sources like CNN generally feature higher credibility compared to self-media posts on Twitter, which may have a higher proportion of true news [11]. This imbalance in the proportion of real and fake news across platforms introduces another classic challenge in domain adaptation: the label shift problem [12]. These two domain shifts underscore the necessity for a benchmark dataset for fake news to evaluate a model’s domain adaptation capabilities.
@@
Figure 1: The proposed FineFake: a multi-domain dataset that encompasses instances from diverse platforms and topics. Each sample is associated with corresponding image, accurate knowledge and fine-grained label.
To alleviate the aforementioned challenges of domain shift, external knowledge graphs like ConceptNet [13] are incorporated to provide extra cross-domain information, serving as a bridge between disparate domains [14, 15, 16]. The determination of the veracity of news often relies on foundational knowledge and latest external information as auxiliary factors for assessment. However, the same entity may have different meanings across domains, causing entity ambiguity. As depicted in Fig 1, the term “apple” within the health domain refers to a fruit, whereas in the business domain, it denotes the company Apple Inc. Consequently, these disparities across domains give rise to the challenge of entity ambiguity, which can introduce noisy knowledge and degrade the performance of the model. Therefore, the development of datasets containing accurate common knowledge is essential for advancing fake news detection.
@@
TABLE II: Comparison between FineFake and other fake news detection datasets.
@@
Breaking![7] 649 BS Detector US Election Three Category
Weibo21[17] 9128 Weibo Nine Topics Real/ Fake
LIAR[6] 12,836 Politifact Political Six Category
Evons[18] 92,969 Media-source US Election Real/ Fake
Weibo[8] 9,528 Weibo Real/ Fake
RD-E[5] 19,162 Snopes/PolitiFact Six Category
Pheme[5] 5,802 Twitter News Events Real/ Fake
FauxBuster[19] 917 Twitter/Reddit Real/ Fake
Twitter[4] 15,629 Twitter Real/ Fake
MM-Covid[20] 11,173 FullFact Health Real/ Fake
MuMIN[21] 984 Twitter Three Category
MR2[22] 14,700 Twitter/Weibo Three Category
@@
FineFake 16,909 Snopes Social Media (e.g. Twitter) Official News (e.g. CNN) Politics, Enter. Business, Health Society, Conflict Real,Text/Image Fake, Text-image/ Content-knowledge Inconsistency, Others
Another significant aspect of fake news across different topics or platforms lies in the diverse underlying strategies used for fabrication. For instance, Twitter has a more proportion of news with fake images than CNN, which can be attributed to the higher prevalence of manipulated images on social media platforms compared to official news sources [23, 24]. Therefore, it is essential to identify distinguishing features of fake news that can provide a rational basis for judgment across platforms. However, conventional fake news detection datasets typically classify news articles into binary classes (real or fake) or employ broad categories such as “most likely” or “somewhat likely” [6, 25]. Unfortunately, such coarse-grained annotations fail to reveal the factors contributing to the falseness of a news item, such as fabricated images or inconsistencies between text and image. Therefore, there’s an urgent need for a fine-grained annotation strategy to uncover the reasons behind fake news.
@@
In order to address the aforementioned challenges, in this paper, we propose a comprehensive and knowledge-enhanced dataset for fake news detection, dubbed FineFake. As illustrated in Table II, FineFake surpasses its predecessors by spanning multiple topics and platforms, enjoying accurate common knowledge and fine-grained annotations, thereby furnishing robust and solid data support for further research endeavors. FineFake encompasses instances collected from diverse media platforms, such as CNN, Reddit and Snopes. All instances are also labeled into six distinct topics (i.e., politics, entertainment, business, health, society and conflict). The inclusion of this comprehensive multi-domain dataset fosters a deeper comprehension of correlations between fake news from various domains. Each news article contains textual content, images, possible social connections, and other pertinent meta-data. To ensure the provision of reliable common knowledge, each news is appended with the relevant knowledge entities and descriptions in a semi-manual labeling manner. Moreover, we introduce an innovative annotation guideline that extends beyond traditional binary class labels. Our approach incorporates a six-category annotation strategy that sheds light on the reasons behind the detected fake news, including real, textual fake, visual fake, text-image inconsistency, content-knowledge inconsistency, and other samples. Based on FineFake, we conduct extensive experiments on data characteristics, fine-grained classification and domain adaptation, establishing a valuable benchmark for future research. Furthermore, to solve the covariate and label shift issues in FineFake, we propose a knowledge-enhanced domain adaptation network, dubbed KEAN, which achieves SOTA performance in most scenarios. Our contributions are summarized as:
@@
  • FineFake dataset represents a pioneering effort for cross-domain fake news detection, which systematically gathers and formalizes multi-modal news content from diverse topics and platforms.
@@
II Related Work
II-A Fake News Detection Datasets
@@
Since the rapid development of the internet, a proliferation of publicly available datasets concerning the detection of fake news has ensued. Initially, researchers predominantly focused on collecting textual data, concentrating on specific domains to construct these datasets [6, 26, 7, 27]. For instance, the LIAR dataset [6] harnesses Politifact [^1] http://www.politifact.com/ to extract news items, subsequently annotating them with six classification labels. Similarly, FEVER [26] is generated by altering sentences extracted from Wikipedia and pre-processing Wikipedia data, each claim is annotated to a three-way classification label. Additionally, the COVID-19 dataset [27] meticulously curates a manually annotated repository of social media posts and articles pertaining to COVID-19, aiming to facilitate research endeavors in identifying pertinent rumors that possess the potential to instigate significant harm.
@@
The aforementioned uni-modal datasets primarily concentrate on linguistic analysis, thus disregarding the crucial dimensions of social networks for dissemination, corresponding images, and metadata, essential for a more comprehensive detection framework. In stark contrast to traditional textual news media releases, the presence of multimodal news incorporating images or videos tends to captivate greater attention and propagate more extensively, thus may lead to more damage transmission. Consequently, there has been a discernible surge in the construction of datasets integrating images and social network data [28, 19, 21, 5, 4]. For instance, MM-COVID [20] offers a multilingual dataset encompassing news articles augmented with pertinent social context and images, aimed at facilitating the detection and mitigation of fake news pertaining to the COVID-19 pandemic. Similarly, Weibo [8] collects original tweet texts, attached images, and contextual information sourced from Weibo, a prominent Chinese microblogging platform renowned for its objective ground-truth labels.
@@
Although the existing multi-modal datasets contain multi-modal data like images, the majority of them overlook critical factors such as multi-domain attributions, external knowledge and fine-grained classification annotations. $MR^{2}$ [22] attempts to address this gap by incorporating evidence retrieved from online sources as metadata to enhance fake news detection. Nonetheless, this approach of online retrieval lacks a guarantee of the accuracy of external evidence, potentially introducing extraneous noise information instead. Additionally, Weibo21  [17] divides textual news data into nine distinct topics, yet it only focus on topics, overlooking the the multitude of platforms through which news dissemination occurs. Although LIAR [6] and RD-E [25] categorize fake news into six categories, they rely on pre-existing labels from fact-checking websites such as “most likely” or “somewhat likely”, failing to explore the underlying reasons behind the identification of fake news.
II-B Fake News Detection Methods
@@
At the outset, a considerable amount of research focused on refining the extraction of semantic features inherent in news content itself, recognizing the wealth of information embedded within the content conducive to discerning its veracity [29, 9, 30, 31]. However, the escalating convergence of semantic structures between fake and authentic news has rendered the task of distinguishing between them based solely on semantics increasingly formidable [1]. Consequently, attention has shifted towards leveraging external knowledge as supplementary information to bolster fake news detection efforts [32, 14, 33, 34, 16]. For instance, CompareNet [16] constructs a directed heterogeneous document graph to compare news to external knowledge base through the extractied entities.
@@
Nonetheless, news content exhibits significant variations across diverse platforms and topics, and the distribution of fake and authentic news also fluctuates accordingly [1, 35]. Effective deployment of a well-trained fake news detection model in real-world scenarios necessitates robust cross-domain capabilities. However, only a limited number of studies have earnestly tackled the challenges posed by multi-domain and cross-domain fake news detection [36, 17, 35, 37]. Consequently, we construct a multi-domain knowledge-enhanced multimodal fine-grained dataset that holds immense potential for facilitating research in the realms of multi-domain and cross-domain fake news detection in reality.
@@
III FineFake: The Knowledge-Enriched Fine-Grained Multi-Domain Dataset
@@
TABLE III: Mappings from the topic categories in Snopes to topics in FineFake.
@@
III-A Multi-domain News Collection
Current fake news detection datasets predominantly gather data from singular platforms within narrow topics [1]. To develop a large-scale dataset spanning multiple platforms and topics with substantial diversity, we employ a comprehensive data collection strategy involving three primary channels: Snopes, social media platforms (e.g., Twitter), and official news websites (e.g., CNN). The overall collection process for FineFake is illustrated in Figure 2.
Figure 2: The Construction Process of FineFake. Snopes is used as the starting point for data collection and the external links within the claim explanations are leveraged as sources for multi-platform data collection. Platforms are categorized into official news platforms and social media platforms while FineFake also collects potential social network information from social media platforms. Finally, each piece of news undergoes semi-manual knowledge annotation and fine-grained annotation to ensure label accuracy.
Snopes  [^2] https://www.snopes.com is a website that verifies the authenticity of news reports. Each verified claim encompasses a substantial amount of additional information regarding the original news post, including external links to the original source, expert-assigned topic labels, and authenticity assessments provided by professionals. The provision of links for news provenance can enrich the sources of news across multiple platforms, while the tagging of topics facilitates the construction of multi-topic databases. Moreover, the authenticity ratings provided by professionals are advantageous for our exploration of the underlying causes of fake news. Unlike previous approaches that only use the summarized claims as samples [22], we collect data on various topics through the topic categorization provided by Snopes, preserving elements such as claims, images, tags, verification materials and rating categories. The external news links are then used as springboards to gather data from multiple platforms for multi-source news collection.
For official news websites, APNews, CNN, New York Times, The Washington Post, and the CDC are chosen due to their credibility and the high quality of content. The tailored web crawlers are programmed to meticulously extract not just the textual content but also images, authors, publication dates and other relevant metadata from each article. In parallel, we extend our data collection framework to integrate social interactions from social media platforms, including Twitter and Reddit, which contain abundant social information that is invaluable for understanding how online news spreads and evolves. For Twitter, the well-documented API is utilized to collect tweets, retweets, replies of the news links found on Snopes. To effectively gather data from Reddit, a specialized crawler is designed to obtain posts and user interactions due to its lack of API, enabling us to build another layer of social network reflecting the news events.
Since data in Snopes is affiliated with tags, to facilitate the study of distinctions among different topics, these tags are consolidated into six categories within the Snopes dataset. The six topic categories, including politics, entertainment, business, health, society and conflict are carefully chosen based on their prevalence and relevance in both Snopes tags and fake news detection [38, 17]. The specific meanings of thses six topics are described as follows:
  • Politics. The politics topic records the statements made by many politicians during elections and political activities, along with corresponding news photographs.
  • Entertainment. The entertainment topic focuses on hot news events and related reports about celebrities in the entertainment industry.
  • Business. The business topic represents news about the business plans and actions of entrepreneurs and well-known companies.
  • Health. The health topic aims at news related to major health events like COVID-19, fake news often leads to social panic and the spread of incorrect treatment methods.
  • Society. The society topic includes news about society events, such as education reform, economic demelopment, residents’ quality of life and so on.
  • Conflict. The conflict topic focuses on news about wars, military and conflict between countries and regions.
The correlation between the category labels in Snopes and our designated topic labels is elucidated in Table III. News collected from the other two platforms are annotated with the same topic as the source data in Snopes.